OpenStudy (anonymous):

Convert 61.13° to degrees, minutes, and seconds. 61°7'48" 61°7'13" 61°7'36" 61°7'54"

OpenStudy (jdoe0001):

well, you have \(61^o\) already

OpenStudy (jdoe0001):

so, how many minutes in 0.13 decimals of a degree?

OpenStudy (jdoe0001):

$$ \cfrac{60'}{1^o}=\cfrac{x'}{0.13^o} \implies x' = \cfrac{60'\times 0.13^o}{1^o} $$

OpenStudy (anonymous):

@jdoe0001 i dont understand above

OpenStudy (anonymous):

Keep multiplying the decimal points by 60 to get 61°7'48"